OpenStudy (anonymous):

4+-1=............. plz help will give medals and become a fan

OpenStudy (anonymous):

Is this a serious question?

OpenStudy (destinymasha):

3

OpenStudy (anonymous):

yes

OpenStudy (destinymasha):

4 + -1 is like subtracting 1 from 4

OpenStudy (anonymous):

4 + (-1) basically becomes 4 -1 so the answer will be 3.

OpenStudy (destinymasha):

so 3

OpenStudy (anonymous):

oh yea ty

OpenStudy (destinymasha):

welcome

OpenStudy (anonymous):

\[ 4+-1=4+0-1=4-1 \]

OpenStudy (davidusa):

hi wio